OpenStudy (anonymous):

Challenge: Graph the equation with a diameter that has endpoints at (-3, 4) and (5, -2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (jdoe0001):

hmm what's the distance between (-3, 4) and (5, -2). ? 

OpenStudy (anonymous):

eq.of circle is (x-x1)(x-x2)+(y-y1)(y-y2)=0

OpenStudy (anonymous):

About 9.90

OpenStudy (anonymous):

So I divide that by two to get the radius right? And with it the center

OpenStudy (jdoe0001):

@sewandowski hmmm how did you get 9.9?

OpenStudy (anonymous):

Square root of: (-3-4)^2 + (5-(-2))^2

OpenStudy (anonymous):

mid point is the center

OpenStudy (anonymous):

Oooh, so I use the midpoint formula?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

hmm anyhow as @surjithayer said find the diameter to get the "radius" ahd the midpoint formula to get the center

OpenStudy (jdoe0001):

btw, the distance aint' 9.9 either

OpenStudy (anonymous):

Ok, so the midpoint of the points is (1,1), which is the center

OpenStudy (jdoe0001):

yeap...1,1 is the center

OpenStudy (jdoe0001):

and the diameter?

OpenStudy (anonymous):

Ok, now I got 10

OpenStudy (anonymous):

so the radius is 5

OpenStudy (jdoe0001):

yeap radius is 5, diameter is 10 then \(\large (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{purple}{ r}}\)

OpenStudy (anonymous):

Perfect, now all I need to do is graph

OpenStudy (anonymous):

Thanks guys!!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

eq. of circle is \[\left\{ x-\left( -3 \right) \right\}\left( x-5 \right)+\left( y-4 \right)\left\{ y-\left( -2 \right) \right\}=0\] solve it.